On January 1, a company issues bonds dated January 1 with a par value of $460,000. The bonds mature in 5 years. The contract rate is 7%, and interest is paid semiannually on June 30 and December 31. The market rate is 8% and the bonds are sold for $441,361. The journal entry to record the first interest payment using straight-line amortization is:

  • (A) debit Interest Expense $17,963.90; credit Premium on Bonds Payable $1,863.90; credit Cash $16,100.00.

  • (B) debit Interest Expense $16,100.00; credit Cash $16,100.00.

  • (C) debit Interest Payable $16,100.00; credit Cash $16,100.00.

  • (D) debit Interest Expense $14,236.10; debit Discount on Bonds Payable $1,863.90; credit Cash $16,100.00.

  • (E) debit Interest Expense $17,963.90; credit Discount on Bonds Payable $1,863.90; credit Cash $16,100.00.
Answer #1
Total Discount on issue of bond = Par value of bonds issued - Issue value of bonds issued = 460000 - 441361 18639
Amortization of discount on issue of bond per year = Total discount on issue of bonds / Maturity period of bonds = 18639 / 5 3727.80
Semi-annual amortization of discount on bonds = Amortization of discount on issue of bond per year * 6/12 = 3727.80 * 6/12 1863.90
Semi-annual interest = Par value of bonds * Contract rate * 6/12 = 460000 * 7% * 6/12 16100.00
Journal entry :
Debit Credit
Interest expense 17963.90
Discount on bonds payable 1863.90
Cash 16100.00
So, the answer is option E